# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ 2.2 Evolution of Each Problem in VLM Era
(a) Sensory AD $\rightarrow$ CLIP-based AD  Sensory AD has continued to develop as a common problem setting for CLIP-based AD, inheriting the challenges of traditional sensory AD [27, 61, 62, 63, 64, 62, 29, 65, 33, 35]. As shown in Table I, the first appearance in a top venue was at CVPR 2023, and since then, a total of six papers have been published in top venues. In addition, there are numerous other papers [63, 61, 64, 66, 65]. Moreover, in terms of benchmarks, in addition to the commonly used MVTec-AD [52], the largest industrial anomaly detection dataset VisA [53] has also become a standard benchmark in the field. Therefore, it is evident that sensory AD has become a highly active and noteworthy field in the VLM era.
@@
(b) Semantic AD/ND $\rightarrow$ Inactive  Research on semantic AD/ND appears to become inactive in the VLM era. As shown in Table I, there are only two papers, TMLR 2022 [37] and CVPR 2024 [35]. However, the CVPR 2024 work [35] aims to build a generalist anomaly detector that solves many AD tasks, including sensory AD and semantic AD, and is not primarily focused on semantic AD. The reasons for the inactivity include saturation of performance for one-class semantic AD/ND, and incompatibility of methods with CLIP for multi-class semantic AD/ND. As for one-class semantic AD/ND, TMLR [37] exists, but the performances with common CIFAR and ImageNet-30 datasets have already achieved around 99%. As for multi-class semantic AD/ND, a common approach is to treat ID classes as a single class, but treating ID classes as a single class is less compatible with CLIP’s class-wise discriminative capability.
@@
(c) OSR $\rightarrow$ CLIP-based OOD Detection  We consider that OSR has been integrated into CLIP-based hard OOD detection. According to Table I, there are no top venue publications on OSR research in the VLM era. Originally, the main difference between OSR and OOD detection was the benchmark setup [22, 55]. OSR typically divides the classes in the one dataset into some known (ID) classes and unknown (OOD) classes, as seen in MNIST-4/6 [67] CIFAR-4/6 [68], CIFAR-50/50 [69], and TinyImageNet-20/180 [70]. However, in recent years, some works on CLIP-based OOD detection incorporate the benchmark setup of OSR and create new benchmarks such as ImageNet-10/ImageNet-20 [26] and ImageNet-protocol [54, 47] for hard OOD detection. Therefore, the boundary between OOD detection and OSR has effectively disappeared, and all research in the VLM era has been integrated into OOD detection.
@@
Nevertheless, while pure OSR research is declining, some studies have used the term “open-set” in the context of domain generalization [71]. These studies deviate from the original scope of OSR research and are rather closely aligned with the field of domain generalization [72]. Therefore, within our generalized OOD detection v2, we do not classify these studies as falling under OSR research. We will discuss them in the context of full-spectrum OOD detection, a research field combining generalization and detection, in Sec. 7.2.
@@
(d) OOD Detection $\rightarrow$ CLIP-based OOD Detection  OOD detection is a highly active research area in the VLM era. As shown in Table I, there are many papers in top venues, indicating a high interest from the community. Additionally, as mentioned above, OSR has been integrated with OOD detection as a field of hard OOD detection [26, 47]. Therefore, it is expected that OOD detection will continue to grow and develop further.
@@
(e) OD $\rightarrow$ Inactive  OD has become less active in the VLM era. Previously, OD was used for open-set semi-supervised learning [73, 74, 75], learning with open-set noisy labels [76], and novelty discovery [77, 78, 79, 80, 81]. The reason for the inactivity is that the use of CLIP led to a reduction in training costs and only a small amount of data needs to be collected, eliminating the need for large amounts of unlabeled data and reducing the need to consider noisy data. However, recently, Liang et al. [49] proposed Unsupervised Universal Fine-Tuning, a new problem setting for CLIP-based OD in ICML2024. Unsupervised Universal Fine-Tuning assumes a more realistic problem setting for unsupervised tuning of the downstream task with CLIP where some OOD samples are included in the unlabeled samples. With this new problem setting, there is still a possibility that OD will become active in the future. However, as OD is not currently an active area, we exclude OD from the main discussion of this survey. Unsupervised Universal Fine-Tuning is deeply related to OOD detection and will be discussed in detail in Sec. 4.3.
@@ 2.3 Discussion
Through Sec. 2.2, we found that previously mixed fields have been correctly organized in the VLM era, and that the focus should be on OOD detection and sensory AD. These fields are still developing, with an increasing number of methodologies and benchmarks, and are expected to become more active in the future. Note here that this does not mean that other fields have come to an end. For example, one reason why one-class semantic AD/ND has not been studied is the saturation of performance [37]. If more fine-grained and challenging datasets could be constructed, the field could be reactive. We put this in out-of-scope for this survey paper, but this is an important future challenge.
@@ 3 Overview of Each Problem in VLM Era
In addition to the above inter-field evolution, we emphasize that the advent of VLMs has significantly changed the field of OOD detection itself. In this section, we present an overview of CLIP-based OOD detection, highlighting the key changes in the problem definition, the problem setting, and benchmarks. In addition, we also present an overview of CLIP-based AD in the hope that the understanding of each field will lead to a deeper understanding of CLIP-based OOD detection. For items that remain unchanged from traditional problems, such as background, applications, and evaluation, we refer the readers to the original generalized OOD detection paper [22].
3.1 CLIP-based Out-of-Distribution Detection
@@
Definition  The definition of CLIP-based OOD detection differs significantly from that of conventional OOD detection. Conventional OOD detection aims to detect test samples drawn from a distribution that is different from the training distribution. As another definition, OOD detection is defined as a task to detect test samples that the model cannot or does not want to generalize [22]. However, for CLIP-based OOD detection, CLIP has a vast amount of knowledge, so the OOD sample is completely unrelated to the distribution of the CLIP’s pretraining data or the CLIP’s own generalization ability. Therefore, traditional definitions cannot adequately describe the definition of CLIP-based OOD detection.
@@
Unlike the previous definition, CLIP-based OOD detection is defined as follows [26, 39]: CLIP-based OOD detection aims to detect samples that do not belong to any ID class text provided by the user. Given a pre-trained model, a classification task of interest is defined by a set of class labels $\mathcal{Y}_{\text{ID}}$, which we refer to as the ID classes. The semantic distribution is represented by the distribution $P(\mathcal{Y}_{\text{ID}})$. CLIP-based OOD detection aims to detect test samples that come from the distribution with the semantic shift from the ID classes, i.e., $P(\mathcal{Y}_{\text{ID}})\neq P(\mathcal{Y}_{\text{OOD}})$. Following the definition of the generalized OOD detection framework [22], ideal OOD detectors should keep the classification performance on test samples from ID class space $\mathcal{Y}_{\text{ID}}$, and reject OOD test samples with semantics outside the support of $\mathcal{Y}_{\text{ID}}$.
@@
Problem Setting  CLIP-based OOD detection focuses on solving the image classification task in a computationally efficient way. Unlike traditional OOD detection settings, which primarily involve training an ID classifier with whole ID data, CLIP-based OOD detection primarily focuses on a zero-shot [26] (i.e., without utilizing ID images) or few-shot [28] (i.e., utilizing only a few ID images) setting. Each detailed definition of both settings are described later in Sec. 4. The field is advancing towards greater computational efficiency, requiring minimal or no training data.
@@
Benchmark  Most recent works in CLIP-based OOD detection use high-resolution and large-scale datasets such as ImageNet [26, 28, 46, 47, 48]. The common ImageNet OOD benchmark uses ImageNet as ID and other datasets [82, 83, 84, 85] as OOD. However, in this common benchmark, the semantics between ID and OOD are far, which may allow easy distinction between the ID and OOD. Therefore, recent works use more challenging OOD benchmarks where they split ImageNet classes into ID and OOD categories for hard OOD detection [26, 47, 86]. The representative datasets are ImageNet-20 [26], ImageNet-10 [26], and the recently proposed ImageNet-protocol [54] created by dividing into multiple variations of ID/OOD pairs from ImageNet-1K. This creation strategy initially focused on OSR but has recently been repurposed for OOD detection. These changes in the datasets shift OOD detection closer to the real world and make it a more challenging and practical task.
3.2 CLIP-based Anomaly Detection
@@
Definition  Unlike OOD detection, the definition of anomaly detection (AD) has not changed between conventional AD and CLIP-based AD. AD is intended for use in specific circumstances (industrial inspection), where samples that deviate from predefined normality are considered an anomaly [22, 11]. Whether a model can generalize is irrelevant to the definition of “Anomaly”. Therefore, even with the emergence of CLIP, the definition has not changed.
@@
Problem Setting  CLIP-based AD focuses on solving anomaly classification and segmentation in a computationally efficient way. Anomaly classification, like conventional AD, is a binary classification task that distinguishes between normality and abnormality. Anomaly segmentation, also following conventional AD, involves segmenting the location of anomalies. Like CLIP-based OOD detection, CLIP-based AD also primarily focuses on a zero-shot [27] (i.e., without utilizing images in the target dataset) or few-shot [27] (i.e., using only a few normal images in the target dataset) setting. Each detailed definition of zero-shot and few-shot settings is described later in Sec. 5. As another shift, conventional AD created separate models for each category  [87, 88, 89, 90, 91, 92, 93], while CLIP-based AD creates a single unified model across multiple categories [27, 61, 29, 63, 35], which leads to a more computationally efficient approach.
@@
One key difference from CLIP-based OOD detection is that CLIP-based OOD detection does not involve localization tasks, while these are mainstream in CLIP-based AD. This will be discussed in detail in Sec. 5.4.
@@
Benchmark  Most works on CLIP-based AD tackle industrial inspection [52, 94, 95]. As for the benchmarks, besides the common MVTec-AD dataset [52], the more challenging dataset VisA [53] has been newly employed [27]. The VisA benchmark includes objects with complex structures such as printed circuit boards and multiple instances with different locations within a single view, making it one of the most challenging datasets currently available in the open datasets. Since the pioneering work in CLIP-based AD (i.e., WinCLIP [27]) used MVTec-AD and VisA, many subsequent works have also used these datasets [63, 33, 65].
@@
Figure 3: Timeline for representative methodologies for CLIP-based out-of-distribution detection and CLIP-based anomaly detection. We observe that an increasing number of methods have recently been proposed for both tasks, indicating the growing activity in these fields.
@@
TABLE II: Representative paper list for CLIP-based out-of-distribution detection and CLIP-based anomaly detection.
@@ Methods
MCM [26], GL-MCM [96], SeTAR [97]
@@
LSN [44], NegPrompt [47], IDPrompt [46]
@@
4 CLIP-based OOD Detection: Methodology
In this section, we introduce the methodologies for CLIP-based out-of-distribution (OOD) detection. Fig. 3 presents the timeline for representative methodologies for CLIP-based OOD detection. Table II presents representative methods. We introduce methods for zero-shot OOD detection in Sec. 4.1, few-shot OOD detection in Sec. 4.2, and other research directions in Sec. 4.3. For each methodology, we categorize them by the type of training and whether additional OOD prompts were employed.

### v2

@@ 2.3 Generalized OOD Detection v2: Evolution in VLM Era
(a) Sensory AD $\rightarrow$ VLM-based AD  Sensory AD has continued to develop as a common problem setting for VLM-based AD, inheriting the challenges of traditional sensory AD (Jeong et al., 2023; Deng et al., 2023; Chen et al., 2023c; b; Tamura, 2023; Chen et al., 2023c; Zhou et al., 2024a; Gu et al., 2024a; Li et al., 2024c; Zhu & Pang, 2024). As shown in Table 1, the first appearance in a top venue was at CVPR 2023, and since then, numerous papers have been published in top venues. Therefore, it is evident that Sensory AD has consistently been an active research field.
@@
(b) Semantic AD/ND $\rightarrow$ Inactive  Research on semantic AD/ND appears to become inactive in the VLM era. As shown in Table 1, there are three papers, TMLR 2022 (Liznerski et al., 2022), CVPR 2024 (Zhu & Pang, 2024), and ICLR 2025 (Yun et al., 2025). However, the CVPR 2024 (Zhu & Pang, 2024) and ICLR 2025 (Yun et al., 2025) work aims to build a generalist anomaly detector that solves many AD tasks, including sensory AD and semantic AD, and is not primarily focused on semantic AD. The reasons for the inactivity include saturation of performance for one-class semantic AD/ND, and incompatibility of CLIP and multi-class semantic AD/ND settings. As for one-class semantic AD/ND, TMLR (Liznerski et al., 2022) exists, but the performances with common CIFAR and ImageNet-30 datasets have already achieved around 99%. As for multi-class semantic AD/ND, a common approach is to treat ID classes as a single class, but treating ID classes as a single class is less compatible with CLIP’s class-wise discriminative capability.
@@
(c) OSR $\rightarrow$ VLM-based OOD Detection  We consider that OSR has been integrated into VLM-based hard OOD detection. According to Table 1, there is only one top venue publication (Miller et al., 2024) on OSR research in the VLM era. Originally, the main difference between OSR and OOD detection was the benchmark setup (Yang et al., 2024; Salehi et al., 2022). OSR typically divides the classes in the one dataset into some known (ID) classes and unknown (OOD) classes, as seen in MNIST-4/6 (Deng, 2012) CIFAR-4/6 (Krizhevsky et al., 2009a), CIFAR-50/50 (Krizhevsky et al., 2009b), and TinyImageNet-20/180 (Torralba et al., 2008). However, in recent years, some works on VLM-based OOD detection incorporate the benchmark setup of OSR and create new benchmarks such as ImageNet-10/ImageNet-20 (Ming et al., 2022a) and ImageNet-protocol (Palechor et al., 2023; Li et al., 2024b) for hard OOD detection. The only work (Miller et al., 2024) discusses the OSR problem in the VLM era, but they do not mention the differences from VLM-based OOD detection, and the definitions of VLM-based OSR are identical to those of VLM-based OOD detection. Considering the number of papers on VLM-based OOD detection and its identical definitions with OSR, we consider that the boundary between OOD detection and OSR has effectively disappeared, and all research in the VLM era has been integrated into OOD detection.
@@
Nevertheless, while pure OSR research is declining, some studies have used the term “open-set” in the context of domain generalization (Shu et al., 2023). These studies deviate from the original scope of OSR research and are rather closely aligned with the field of domain generalization (Zhou et al., 2022a). Therefore, within our generalized OOD detection v2, we do not classify these studies as falling under OSR research.
@@
(d) OOD Detection $\rightarrow$ VLM-based OOD Detection  OOD detection is a highly active research area in the VLM era. As shown in Table 1, there are many papers in top venues, indicating a high interest from the community. Additionally, as mentioned above, OSR has been integrated with OOD detection as a field of hard OOD detection (Ming et al., 2022a; Li et al., 2024b). Therefore, it is expected that OOD detection will continue to grow and develop further.
@@
(e) OD $\rightarrow$ Inactive  OD has become less active in the VLM era. Previously, OD was used for open-set semi-supervised learning (Yu et al., 2020; Saito et al., 2021; Cao et al., 2022), learning with open-set noisy labels (Wang et al., 2018), and novelty discovery (Han et al., 2019; Zhao & Han, 2021; Jia et al., 2021; Vaze et al., 2022b; Joseph et al., 2022). The reason for the inactivity is that the use of CLIP led to a reduction in training costs and only a small amount of data needs to be collected, eliminating the need for large amounts of unlabeled data and reducing the need to consider noisy data. However, recently, Liang et al. (2024) proposed Unsupervised Universal Fine-Tuning, a new problem setting for VLM-based OD in ICML2024. Unsupervised Universal Fine-Tuning assumes a more realistic problem setting for unsupervised tuning of the downstream task with CLIP where some OOD samples are included in the unlabeled samples. With this new problem setting, there is still a possibility that OD will become active in the future. However, as OD is not currently an active area, we exclude OD from the main discussion of this survey. Unsupervised Universal Fine-Tuning is deeply related to OOD detection and will be discussed in detail in Sec. 4.3.
@@ 2.4 Discussion
Through Sec. 2.3, we found that previously mixed fields have been correctly organized in the VLM era, and that the focus becomes OOD detection and sensory AD. These fields are still developing, with an increasing number of methodologies and benchmarks, and are expected to become more active in the future. Note here that this does not mean that other fields have come to an end. For example, one reason why one-class semantic AD/ND has not been studied is the saturation of performance (Liznerski et al., 2022). If more fine-grained and challenging datasets could be constructed, the field could be reactive. We put this in out-of-scope for this survey paper, but this is an important future challenge.
@@ 3 Overview of Each Problem in VLM Era
In addition to the above inter-field evolution, we emphasize that the advent of VLMs has significantly changed the field of OOD detection itself. In this section, we present an overview of VLM-based OOD detection, highlighting the key changes in the problem definition, the problem setting, and benchmarks. In addition, we also present an overview of VLM-based AD in the hope that the understanding of each field will lead to a deeper understanding of VLM-based OOD detection. We describe the changes in problem definition, problem setting, and benchmarks in the VLM era. As for the background, applications, and evaluation in each field, we refer the readers to the original generalized OOD detection paper (Yang et al., 2024).
3.1 VLM-based Out-of-Distribution Detection
@@
Definition  The definition of VLM-based OOD detection differs significantly from that of conventional OOD detection. Conventional OOD detection aims to detect test samples drawn from a distribution that is different from the training distribution. As another definition, OOD detection is defined as a task to detect test samples that the model cannot or does not want to generalize (Yang et al., 2024). However, for VLM-based OOD detection, CLIP has a vast amount of knowledge, so the OOD sample is completely unrelated to the distribution of the CLIP’s pretraining data or the CLIP’s own generalization ability. Therefore, traditional definitions cannot adequately describe the definition of VLM-based OOD detection.
@@
Unlike the previous definition, VLM-based OOD detection is defined as follows (Ming et al., 2022a; Esmaeilpour et al., 2022): VLM-based OOD detection aims to detect samples that do not belong to any ID class text provided by the user. Given a pre-trained model, a classification task of interest is defined by a set of class labels $\mathcal{Y}_{\text{ID}}$, which we refer to as the ID classes. The semantic distribution is represented by the distribution $P(\mathcal{Y}_{\text{ID}})$. VLM-based OOD detection aims to detect test samples that come from the distribution with the semantic shift from the ID classes, i.e., $P(\mathcal{Y}_{\text{ID}})\neq P(\mathcal{Y}_{\text{OOD}})$. Following the definition of the generalized OOD detection framework (Yang et al., 2024), ideal OOD detectors should keep the classification performance on test samples from ID class space $\mathcal{Y}_{\text{ID}}$, and reject OOD test samples with semantics outside the support of $\mathcal{Y}_{\text{ID}}$.
@@
Problem Setting  VLM-based OOD detection focuses on solving the image classification task in a computationally efficient way. Unlike traditional OOD detection settings, which primarily involve training an ID classifier with whole ID data, VLM-based OOD detection primarily focuses on a zero-shot (Ming et al., 2022a) (i.e., without utilizing ID images) or few-shot (Miyai et al., 2023b) (i.e., utilizing only a few ID images) setting. Each detailed definition of both settings is described later in Sec. 4. The field is advancing towards greater computational efficiency, requiring minimal or no training data.
@@
Benchmark  The benchmark has also changed between VLM-based methods and previous approaches. Earlier works before CLIP often utilized small-scale datasets such as CIFAR (Krizhevsky et al., 2009a; b) and MNIST (Deng, 2012). In contrast, most recent works in VLM-based OOD detection use high-resolution and large-scale datasets such as ImageNet (Ming et al., 2022a; Miyai et al., 2023b; Bai et al., 2024a; Li et al., 2024b; Cao et al., 2024a). The common ImageNet OOD benchmark uses ImageNet as ID and other datasets (Van Horn et al., 2018; Zhou et al., 2017; Xiao et al., 2010; Cimpoi et al., 2014) as OOD. However, in this common benchmark, the semantics between ID and OOD are far, which may allow easy distinction between the ID and OOD. Therefore, recent works use more challenging OOD benchmarks where they split ImageNet classes into ID and OOD categories for hard OOD detection (Ming et al., 2022a; Li et al., 2024b; Jung et al., 2024). The representative datasets are ImageNet-20 (Ming et al., 2022a), ImageNet-10 (Ming et al., 2022a), and ImageNet-protocol (Palechor et al., 2023) created by dividing into multiple variations of ID/OOD pairs from ImageNet-1K. This creation strategy initially focused on OSR but has recently been repurposed for OOD detection. More recently,  Noda et al. (2025) proposed ImageNet-X for hard OOD detection and ImageNet-FS-X and Wilds-FS-X for hard full-spectrum OOD detection, which include more challenging conditions reflecting real-world scenarios. These changes in the datasets shift OOD detection closer to the real world and make it a more challenging and practical task.
3.2 VLM-based Anomaly Detection
@@
Definition  Unlike OOD detection, the definition of anomaly detection (AD) has not changed between conventional AD and VLM-based AD. AD is intended for use in specific circumstances (industrial inspection), where samples that deviate from predefined normality are considered an anomaly (Yang et al., 2024; Ruff et al., 2021). Whether a model can generalize is irrelevant to the definition of “Anomaly”. Therefore, even with the emergence of CLIP, the definition has not changed.
@@
Problem Setting  VLM-based AD focuses on solving anomaly classification and segmentation in a computationally efficient way. Anomaly classification is a binary classification task that distinguishes between normality and abnormality. Anomaly segmentation involves segmenting the location of anomalies. Like VLM-based OOD detection, VLM-based AD also primarily focuses on a zero-shot (Jeong et al., 2023) (i.e., without utilizing images in the target dataset) or few-shot (Jeong et al., 2023) (i.e., using only a few images in the target dataset) setting. Each detailed definition of zero-shot and few-shot settings is described later in Sec. 5. As another shift, conventional AD created separate models for each category  (Bergmann et al., 2020; Defard et al., 2021; Li et al., 2021; Liznerski et al., 2021; Yi & Yoon, 2020; Zavrtanik et al., 2021; Thulasidasan et al., 2021), while VLM-based AD creates a single unified model across multiple categories (Jeong et al., 2023; Deng et al., 2023; Zhou et al., 2024a; Chen et al., 2023b; Zhu & Pang, 2024), which leads to a more computationally efficient approach.
@@
One key difference from VLM-based OOD detection is that VLM-based OOD detection does not involve localization tasks, while these are mainstream in VLM-based AD. This will be discussed in detail in Sec. 5.4.
@@
Benchmark  Most works on VLM-based AD tackle industrial inspection (Bergmann et al., 2019; Chu & Kitani, 2020; Atha & Jahanshahi, 2018). As for the benchmarks, MVTec-AD dataset (Bergmann et al., 2019) and VisA dataset (Zou et al., 2022) are commonly used. MVTEC-AD includes 15 classes including 5 different texture categories and 10 different object categories with totally 5,354 high-resolution images. VisA covers 12 objects such as in 3 domain consisting of 10,821 high-resolution color images. The VisA benchmark includes objects with complex structures such as printed circuit boards and multiple instances with different locations within a single view, making it one of the most challenging datasets currently available in the open datasets. As for other datasets, KSDD (Tabernik et al., 2020), BTAD (Mishra et al., 2021), and MPDD (Jezek et al., 2021) are often used. However, apart from MVTec-AD and VisA, the datasets used vary depending on the papers (Yang et al., 2025; Cao et al., 2024c; Qu et al., 2024; Li et al., 2024f).
@@
Figure 3: Timeline for representative methodologies for VLM-based out-of-distribution detection and VLM-based anomaly detection. We observe that an increasing number of methods have recently been proposed for both tasks, indicating the growing activity in these fields.
@@
Table 2: Representative paper list for VLM-based out-of-distribution detection and anomaly detection.
@@ Methods
MCM (Ming et al., 2022a), GL-MCM (Miyai et al., 2025b), SeTAR (Li et al., 2024e), TAG (Liu & Christopher, 2024), CoVer (Zhang et al., 2024a)
@@
LSN (Nie et al., 2023), NegPrompt (Li et al., 2024b), ID-like-Prompt (Bai et al., 2024a), Local-Prompt (Zeng et al., 2025)
@@
4 VLM-based OOD Detection: Methodology
In this section, we introduce the methodologies for VLM-based out-of-distribution (OOD) detection. Fig. 3 presents the timeline for representative methodologies for VLM-based OOD detection. Table 2 presents representative methods. We introduce methods for zero-shot OOD detection in Sec. 4.1, few-shot OOD detection in Sec. 4.2, and other research directions in Sec. 4.3. We categorize each methodology by the training type and the use of OOD prompts. As for the OOD prompts, the methods that employ additional OOD prompts and do not use them are shown in Fig. 4 (a) and (b), respectively. As for the training type, we categorize the methods into training-free, ID training, and auxiliary training (training with other data than ID training data). The illustration of each training scenario is shown in Fig. 5.
Figure 4: Illustration of the OOD detection methods with OOD prompt and without OOD prompt.
Figure 5: Illustration of the OOD detection methods for each training type in both zero-shot and few-shot settings.